# Beyond Single-Turn: A Survey on Multi-Turn Interactions with Large Language Models

### v4

@@ 3.2.3 Knowledge Graph Integration
Many research in this field focuses on deriving graph-structured embeddings (typically using GNN), which are then integrated during the continuous pre-training or fine-tuning stages of LLMs [212, 213, 214]. Some of them utilize existing KGs for obtaining commonsense or domain knowledge that is lacking between conversations. Wang et al. (2024) [212] address the rational response candidate selection problem, where commonsense—often necessary—is explicitly omitted during human-LLM conversational interactions. The system builds KGs based on external commonsense sources. Both the GNN and LLM are jointly fine-tuned using a response selection loss, which measures how effectively the model ranks the correct response among several candidates. Jain et al. (2024) [213] construct a dynamic graph representation of both the ongoing conversation and Wikidata KG (with text, tables, and infoboxes) [249]. The method jointly trains GNN and LLM with additional memory module, helping LLM to maintain context across multiple dialogue turns and enhances its reasoning capabilities for QA answering from heterogeneous sources. In the healthcare domain, Gao et al. (2025) [214] integrates LLMs with Unified Medical Language System–based KGs [250] using a graph isomorphism network to rank and identify knowledge pathways relevant to the clinical contexts of patients, thus enhancing diagnostic processes.
Being able to refer entities and coreferences from past dialogues and conversations can help LLMs with reasoning and instruction following abilities. Consequently, several research initiatives are now focusing on constructing or modifying KGs to integrate historical dialogue data. This integration helps to create richer, context-aware models that not only understand isolated queries but also leverage prior conversational context for improved performance. The works of Wang et al. (2024)[212] and Jain et al. (2024)[213], as discussed above, append existing KGs with entities and relationships from past conversations. SURGE [215] builds KGs that specifically encode relevant knowledge for ongoing conversation, with triplets consisting of entities and their relations as items. The GNN is incorporated to perform multi-hop reasoning and connect disparate pieces of information. Tan et al. (2025) [216] propose to prune existing KGs to remove irrelevant information, incorporating improved graph structures with additional prompting and LLMs to find candidate paths in multi-hop and multi-entity QAs.
Besides directly intergating graphs into training or finetuning LLMs, several research focus on Graph RAG for enhancing reasoning and flexibility in dynamic conversations. Typically, Graph RAG extracts entities and their relationships from documents or dialogues to build a KG. Then, it traverses the knowledge graph to retrieve subgraphs. The retrieved graph information is then integrated into the LLM’s prompt, which allows LLM to generate a response that is richer, more coherent, and better grounded in factual knowledge [219, 217, 218].
@@ 3.3.1 Single Agent Approaches
One influential example is the ReAct by Yao et al. (2023) [251], which integrates explicit reasoning steps with action executions, enabling the agent to interact dynamically with knowledge bases or external APIs. In this paradigm, the agent proactively decides when to retrieve external information, significantly reducing hallucination issues in open-domain question answering and interactive decision-making tasks. Empirical evaluations showed notable performance gains on benchmarks such as HotpotQA [252], FEVER [253], ALFWorld [254], and WebShop [255] compared to static prompting or reinforcement learning baselines.
@@
Extending this concept, the Toolformer framework [221] trains an LLM to autonomously determine the necessity and timing of external tool usage, such as calculators or web search APIs, by inserting specialized API-call tokens within its generations. By learning to integrate tool use in a self-supervised manner, Toolformer achieves substantial accuracy improvements on zero-shot arithmetic, knowledge retrieval, and translation tasks. This method demonstrates how explicit training for proactive tool invocation can significantly enhance multi-turn problem-solving without increasing model size. Similarly, HuggingGPT  [222] employs an LLM as a centralized orchestrator that autonomously decomposes complex user requests into manageable sub-tasks delegated to specialized external models. Although HuggingGPT primarily emphasizes multi-modal and multi-model coordination, its relevance here lies in showcasing an LLM’s capability for sophisticated planning and iterative decision-making across multi-step interactions, reinforcing the power of agent-based decomposition strategies.
@@
The Reflexion framework introduced by Shinn et al. (2023) [223] further pushes the concept of single-agent improvement by enabling self-reflective feedback loops. Instead of relying on traditional gradient-based learning methods, Reflexion employs verbal reinforcement learning, allowing the LLM to record textual reflections of its previous mistakes into episodic memory. Subsequent interactions utilize these reflections for iterative self-improvement, significantly boosting the agent’s performance on coding and sequential decision-making benchmarks, surpassing even highly advanced models like GPT-4 in single-pass accuracy on code-generation tasks such as HumanEval. Extending this iterative self-improvement paradigm, the Voyager agent by Wang et al. (2023) [224] exemplifies lifelong learning capabilities within an open-ended virtual environment (Minecraft). Voyager autonomously engages in continuous loops of planning, code generation, environment-based execution, and observation, incrementally refining its skill repository through persistent, multi-turn interactions. This approach demonstrates impressive exploration efficiency, rapid skill generalization, and improved cumulative task success compared to prior single-agent baselines, highlighting the effectiveness of iterative, experience-driven knowledge acquisition in complex, open-ended environments.
@@
For comprehensive agent evaluation, Liu et al. (2023) [256] introduced AgentBench, a rigorous benchmark designed to assess LLMs on agentic tasks—contexts requiring LLMs to make decisions and execute actions within interactive environments to accomplish specific goals. AgentBench features eight diverse simulated environments, spanning embodied navigation, interactive games, tool utilization, reasoning puzzles, and web interaction, enabling systematic evaluation of LLMs’ reasoning, planning, and decision-making capabilities across extended interaction sequences. The authors conducted an extensive evaluation of 27 LLMs, including both open-source and proprietary API-based models, revealing substantial performance differentials. Their findings demonstrate that leading proprietary models such as GPT-4 and Claude exhibit remarkable proficiency in functioning as coherent agents in complex, long-horizon tasks, often successfully completing scenarios that confound other models. Nevertheless, even these advanced systems demonstrate considerable limitations, with a significant performance gap persisting between them and the most capable open-source alternatives.
@@ 3.3.2 Multi-Agents Approaches
Role-based frameworks assign distinct roles to agents, guiding structured multi-turn dialogues. CAMEL [225] by Li et al. (2023) uses inception prompting to enable autonomous role-playing between agents (e.g., user and assistant), successfully generating conversational data and revealing emergent cognitive behaviors. Extending structured cooperation, ChatDev [226] by Qian et al. (2024) organizes multiple specialized agents into virtual software-development teams (architect, coder, tester) that sequentially interact, significantly improving software quality over single-agent baselines such as GPT-4. Similarly, Dong et al. (2024) [227] proposed structured collaboration among analyst, coder, and tester agents, greatly enhancing code-generation accuracy. Further emphasizing clear workflows, MetaGPT [228] encodes human-like Standard Operating Procedures into prompts, ensuring systematic verification and substantially reducing cascading errors, thus outperforming simpler multi-agent setups. Similarly, the Mixture-of-Search-Agents (MoSA) approach [257] leverages multiple LLMs that propose and refine solutions in tandem. Each agent can independently suggest a next step or critique another agent’s partial solution, and through this multi-turn collaboration the group avoids single-model blind spots. MoSA demonstrated higher accuracy on challenging math sets (e.g. a MATH benchmark subset) than any single model working alone.
Chen et al. (2024) [258] propose BUTTON ("Bottom-Up then Top-Down"), a systematic method to train LLMs for executing multi-step tool use or function calls in conversational contexts. In the bottom-up phase, BUTTON generates simple atomic instruction–function pairs derived from real-world scenarios. The top-down phase creates a simulated environment involving interactions among user agents, assistant agents, and tool agents, emulating realistic multi-turn dialogues. Using this approach, the authors develop BUTTONInstruct, a dataset comprising 8,000 multi-turn dialogues with compositional function-calling tasks. Models fine-tuned on BUTTONInstruct exhibit significant improvements in planning and correctly executing complex sequences of API calls.
@@
Debate-based methods enhance reasoning accuracy by structuring iterative critiques among agents. Du et al. (2023) [229] showed multi-agent debate significantly improves factual correctness and logical reasoning, outperforming single-agent solutions on mathematical and strategic tasks through structured back-and-forth critique. Complementarily, Generative Agents by [259] explore multi-turn social simulations, demonstrating emergent realistic behaviors (planning, social interactions) that arise naturally through iterative dialogue, highlighting the broader implications of structured deliberation in multi-agent interactions.
@@
Dynamic agent-composition methods create flexible agent teams tailored to specific tasks. AutoAgents [230] automatically generates specialized agents, accompanied by an observer agent that monitors and adjusts the interaction dynamically, surpassing fixed-role approaches on heterogeneous tasks. Similarly, AgentVerse [231] provides a versatile platform for agents to dynamically join or leave teams, enhancing adaptability and performance, while simultaneously producing beneficial emergent behaviors like negotiation and consensus formation.
@@
Despite their promising results, multi-agent LLM systems still exhibit notable challenges and limitations. Recent critiques commonly highlight recurring issues such as role misassignments [260, 261], where agents frequently misunderstand or deviate from their assigned responsibilities, causing redundancy, confusion, or omission of crucial tasks. Inefficient communication overhead and compounded errors [260, 261, 262] are also prevalent, stemming from extensive interactions required for coordination among agents, which not only increases computational costs but also propagates and amplifies mistakes made by individual agents. Additionally, inadequate verification mechanisms [260, 261] lead to unreliable or inconsistent outcomes, as agents often fail to effectively validate intermediate or final results, causing errors to remain undetected or improperly resolved. Emergent risks such as miscoordination, conflicts, and unintended collusion among autonomous agents [262] further complicate multi-agent interactions, highlighting deeper issues related to the unpredictability and complexity inherent in systems of multiple interacting entities. Collectively, these insights underline the necessity for future research to prioritize robust coordination protocols, scalable memory-sharing mechanisms, adaptive verification strategies, and comprehensive evaluation frameworks to effectively harness the full potential of multi-agent collaborative intelligence, enhancing multi-turn interactions and paving the way toward robust, adaptive, and human-aligned AI collaboration.
@@ 4.1.1 Context Retention & Coherence
LLMs struggle to maintain long-term context, leading to incoherence or contradictions in extended dialogues. As conversations grow, models often forget or confuse earlier details, causing lapses in consistency. Recent evaluations show that increasing the distance between a query and its relevant prior context degrades model performance [27]. Even advanced chat-oriented models exhibit only modest gains in multi-turn coherence despite larger context windows and alignment tuning [30]. Multi-turn challenge benchmarks reveal that instruction retention and self-coherence remain difficult for current models – they frequently fail to remember instructions or maintain a consistent narrative over several turns [72]. This indicates that preserving conversational state across turns is an open problem.
@@ 4.1.2 Anaphora & Ellipsis Resolution
Multi-turn dialogue requires resolving pronouns, omissions, and references to earlier utterances. Today’s LLMs can easily misinterpret sentences like “That one looks good” or “I did it,” especially in complex dialogues with many entities. Models often falter at linking such utterances to the correct antecedents in prior context. For example, keeping track of characters or user preferences mentioned only implicitly can lead to confusion in later turns. Recent dialogue evaluations emphasize inference memory of user-provided information as a key gap: models frequently forget or mix up attributes (e.g. a user’s name or previously stated facts) when referred to later [72]. Robust anaphora and ellipsis resolution, akin to co-reference resolution in dialogue, remains an open challenge for maintaining coherence across turns.
@@ 4.1.3 Ambiguity Recognition & Clarification
When user inputs are ambiguous or underspecified, aligned LLMs tend to either over-hedge (give vague answers) or implicitly guess the user’s intent, rather than ask clarifying questions. This is a fundamental limitation in context understanding. Chen et al. (2024) [192] report that current conversational agents often do not adequately disambiguate – if faced with an unclear request, they guess or provide a generic response instead of seeking clarification. For instance, an instruction like “Tell me about that report” might prompt an arbitrary guess about which report, rather than a question to identify the reference. The lack of proactive clarification leads to misunderstandings that compound over a dialogue. Developing LLMs that can recognize ambiguity and ask targeted follow-up questions (as humans do) is an open research direction.
@@ 4.2.1 Error Propagation & Compounding
Errors or misunderstandings made by the model (or user) in one turn can be carried into subsequent turns, often amplifying into larger reasoning failures. If an LLM answers a question incorrectly or the user introduces a false premise, the misinformation may persist through the dialogue. Studies have found that most LLMs perform worse in multi-turn settings than in single-turn, partly due to sensitivity to dialogue history – once a mistake enters the context, the model is likely to build on it in later reasoning [27]. This compounding error effect has been noted in sequential decision tasks as well, where minor mistakes accumulate over interaction trajectories. In other words, the model lacks a robust mechanism to correct or “forget” errors mid-dialogue. This challenge is pronounced when the model misinterprets a user question early on or hallucinates a fact – without intervention, subsequent turns will continue down the wrong path. Designing models that can detect and self-correct such errors (or accept user corrections) remains an open problem.

### v5

@@ 3.2.3 Knowledge Graph Integration
Much research in this area focuses on deriving graph-structured embeddings via GNNs and integrating them during continuous pre-training or fine-tuning of LLMs (Wang et al., 2024i; Jain and Lapata, 2024; Gao et al., 2025a). Several works draw on existing knowledge graphs (KGs) to supply commonsense or domain knowledge absent from conversational context. Wang et al. (2024i) tackle rational response candidate selection, where commonsense knowledge is often implicitly omitted in human–LLM interactions; they construct KGs from external commonsense sources and jointly fine-tune GNN and LLM using a response selection loss that measures correct-response ranking. Jain and Lapata (2024) build a dynamic graph over both the ongoing conversation and Wikidata (Waagmeester et al., 2020), jointly training a GNN and LLM with an additional memory module to maintain context across turns and improve reasoning over heterogeneous sources. In healthcare, Gao et al. (2025a) integrate LLMs with UMLS-based KGs (Bodenreider, 2004) via a graph isomorphism network to rank clinically relevant knowledge pathways, enhancing diagnostic reasoning.
Being able to refer entities and coreferences from past dialogues and conversations can help LLMs with reasoning and instruction following abilities. Consequently, several research initiatives are now focusing on constructing or modifying KGs to integrate historical dialogue data. This integration helps to create richer, context-aware models that not only understand isolated queries but also leverage prior conversational context for improved performance. The works of Wang et al. (2024i) and Jain and Lapata (2024), as discussed above, append existing KGs with entities and relationships from past conversations. SURGE (Kang et al., 2023) builds KGs that specifically encode relevant knowledge for ongoing conversation, with triplets consisting of entities and their relations as items. The GNN is incorporated to perform multi-hop reasoning and connect disparate pieces of information. Tan et al. (2025) propose to prune existing KGs to remove irrelevant information, incorporating improved graph structures with additional prompting and LLMs to find candidate paths in multi-hop and multi-entity QAs.
Besides directly intergating graphs into training or finetuning LLMs, several research focus on Graph RAG for enhancing reasoning and flexibility in dynamic conversations. Typically, Graph RAG extracts entities and their relationships from documents or dialogues to build a KG. Then, it traverses the knowledge graph to retrieve subgraphs. The retrieved graph information is then integrated into the LLM’s prompt, which allows LLM to generate a response that is richer, more coherent, and better grounded in factual knowledge (Wang et al., 2024h; Mavromatis and Karypis, 2025; Yang et al., 2025c).
@@ 3.3.1 Single Agent Approaches
One influential example is the ReAct by Yao et al. (2023), which integrates explicit reasoning steps with action executions, enabling the agent to interact dynamically with knowledge bases or external APIs. In this paradigm, the agent proactively decides when to retrieve external information, significantly reducing hallucination issues in open-domain question answering and interactive decision-making tasks. Empirical evaluations showed notable performance gains on benchmarks such as HotpotQA (Yang et al., 2018), FEVER (Thorne et al., 2018), ALFWorld (Shridhar et al., 2021), and WebShop (Yao et al., 2022) compared to static prompting or reinforcement learning baselines.
@@
Extending this concept, the Toolformer framework (Schick et al., 2023) trains an LLM to autonomously determine the necessity and timing of external tool usage, such as calculators or web search APIs, by inserting specialized API-call tokens within its generations. By learning to integrate tool use in a self-supervised manner, Toolformer achieves substantial accuracy improvements on zero-shot arithmetic, knowledge retrieval, and translation tasks. This method demonstrates how explicit training for proactive tool invocation can significantly enhance multi-turn problem-solving without increasing model size. Similarly, HuggingGPT  (Shen et al., 2023b) employs an LLM as a centralized orchestrator that autonomously decomposes complex user requests into manageable sub-tasks delegated to specialized external models. Although HuggingGPT primarily emphasizes multi-modal and multi-model coordination, its relevance here lies in showcasing an LLM’s capability for sophisticated planning and iterative decision-making across multi-step interactions, reinforcing the power of agent-based decomposition strategies.
@@
The Reflexion framework introduced by Shinn et al. (2023) further pushes the concept of single-agent improvement by enabling self-reflective feedback loops. Instead of relying on traditional gradient-based learning methods, Reflexion employs verbal reinforcement learning, allowing the LLM to record textual reflections of its previous mistakes into episodic memory. Subsequent interactions utilize these reflections for iterative self-improvement, significantly boosting the agent’s performance on coding and sequential decision-making benchmarks, surpassing even highly advanced models like GPT-4 in single-pass accuracy on code-generation tasks such as HumanEval. Extending this iterative self-improvement paradigm, the Voyager agent by Wang et al. (2024b) exemplifies lifelong learning capabilities within an open-ended virtual environment (Minecraft). Voyager autonomously engages in continuous loops of planning, code generation, environment-based execution, and observation, incrementally refining its skill repository through persistent, multi-turn interactions. This approach demonstrates impressive exploration efficiency, rapid skill generalization, and improved cumulative task success compared to prior single-agent baselines, highlighting the effectiveness of iterative, experience-driven knowledge acquisition in complex, open-ended environments.
@@
For comprehensive agent evaluation, Liu et al. (2024d) introduce AgentBench, a rigorous benchmark designed to assess LLMs on agentic tasks—contexts requiring LLMs to make decisions and execute actions within interactive environments to accomplish specific goals. AgentBench features eight diverse simulated environments, spanning embodied navigation, interactive games, tool utilization, reasoning puzzles, and web interaction, enabling systematic evaluation of LLMs’ reasoning, planning, and decision-making capabilities across extended interaction sequences. The authors evaluate 27 LLMs, spanning open-source and proprietary API-based models, and find large performance differences. Leading proprietary systems such as GPT-4 and Claude are notably effective as coherent agents on complex, long-horizon tasks, often completing scenarios that defeat other models. However, even these models have clear limitations, and a substantial gap remains between them and the strongest open-source alternatives.
@@ 3.3.2 Multi-Agents Approaches
Role-based frameworks assign distinct roles to agents, guiding structured multi-turn dialogues. CAMEL (Li et al., 2023c) uses inception prompting to enable autonomous role-playing between agents (e.g., user and assistant), successfully generating conversational data and revealing emergent cognitive behaviors. Extending structured cooperation, ChatDev (Qian et al., 2024) organizes multiple specialized agents into virtual software-development teams (architect, coder, tester) that sequentially interact, significantly improving software quality over single-agent baselines such as GPT-4. Similarly, Dong et al. (2024b) propose structured collaboration among analyst, coder, and tester agents, greatly enhancing code-generation accuracy. Further emphasizing clear workflows, MetaGPT (Hong et al., 2024) encodes human-like Standard Operating Procedures into prompts, ensuring systematic verification and substantially reducing cascading errors, thus outperforming simpler multi-agent setups. Similarly, the Mixture-of-Search-Agents (MoSA) approach (Yang et al., 2025a) leverages multiple LLMs that propose and refine solutions in tandem. Each agent can independently suggest a next step or critique another agent’s partial solution, and through this multi-turn collaboration the group avoids single-model blind spots. MoSA demonstrated higher accuracy on challenging math sets (e.g. a MATH benchmark subset) than any single model working alone.
Chen et al. (2025b) propose BUTTON ("Bottom-Up then Top-Down"), a systematic method for training LLMs on multi-step tool use in conversational settings. The bottom-up phase generates atomic instruction–function pairs from real-world scenarios, while the top-down phase simulates multi-turn dialogues among user, assistant, and tool agents. This process yields BUTTONInstruct, a dataset of 8,000 multi-turn dialogues with compositional function-calling tasks. Models fine-tuned on BUTTONInstruct show significant improvements in planning and executing complex API call sequences.
@@
Debate-based methods enhance reasoning accuracy by structuring iterative critiques among agents. Du et al. (2024) show multi-agent debate significantly improves factual correctness and logical reasoning, outperforming single-agent solutions on mathematical and strategic tasks through structured back-and-forth critique. Complementarily, Generative Agents by Park et al. (2023) explore multi-turn social simulations, demonstrating emergent realistic behaviors (planning, social interactions) that arise naturally through iterative dialogue, highlighting the broader implications of structured deliberation in multi-agent interactions.
@@
Dynamic agent-composition methods create flexible agent teams tailored to specific tasks. AutoAgents (Chen et al., 2024a) automatically generates specialized agents, accompanied by an observer agent that monitors and adjusts the interaction dynamically, surpassing fixed-role approaches on heterogeneous tasks. Similarly, AgentVerse (Chen et al., 2024h) provides a versatile platform for agents to dynamically join or leave teams, enhancing adaptability and performance, while simultaneously producing beneficial emergent behaviors like negotiation and consensus formation.
@@
Despite their promising results, multi-agent LLM systems still exhibit notable challenges and limitations. Recent critiques commonly highlight recurring issues such as role misassignments (Han et al., 2024; Cemri et al., 2025), where agents frequently misunderstand or deviate from their assigned responsibilities, causing redundancy, confusion, or omission of crucial tasks. Inefficient communication overhead and compounded errors (Han et al., 2024; Cemri et al., 2025; Hammond et al., 2025) are also prevalent, stemming from extensive interactions required for coordination among agents, which not only increases computational costs but also propagates and amplifies mistakes made by individual agents. Additionally, inadequate verification mechanisms (Han et al., 2024; Cemri et al., 2025) lead to unreliable or inconsistent outcomes, as agents often fail to effectively validate intermediate or final results, causing errors to remain undetected or improperly resolved. Emergent risks such as miscoordination, conflicts, and unintended collusion among autonomous agents (Hammond et al., 2025) further complicate multi-agent interactions, highlighting deeper issues related to the unpredictability and complexity inherent in systems of multiple interacting entities. Collectively, these insights underline the necessity for future research to prioritize robust coordination protocols, scalable memory-sharing mechanisms, adaptive verification strategies, and comprehensive evaluation frameworks to effectively harness the full potential of multi-agent collaborative intelligence, enhancing multi-turn interactions and paving the way toward robust, adaptive, and human-aligned AI collaboration.
@@ 4.1.1 Context Retention & Coherence
LLMs struggle to maintain long-term context across extended dialogues, leading to incoherence and contradictions. As conversations grow, models tend to forget or confuse earlier details: performance degrades as the distance between a query and its relevant prior context increases (Kwan et al., 2024), and even advanced chat-oriented models show only modest gains in multi-turn coherence despite larger context windows and alignment tuning (Bai et al., 2024). Multi-turn benchmarks further reveal that instruction retention and self-consistency remain difficult—models frequently fail to remember instructions or maintain a coherent narrative over several turns (Deshpande et al., 2025). Preserving conversational state across turns thus remains an open problem.
@@ 4.1.2 Anaphora & Ellipsis Resolution
Multi-turn dialogue requires resolving pronouns, omissions, and references to earlier utterances. LLMs routinely misinterpret expressions such as "That one looks good" or "I did it," particularly in complex dialogues with many entities, failing to link utterances to the correct antecedents. Recent evaluations highlight inference memory as a key gap: models frequently forget or conflate user-provided attributes—such as a name or previously stated fact—when referenced implicitly in later turns (Deshpande et al., 2025). Robust anaphora and ellipsis resolution, analogous to coreference resolution in dialogue, remains an open challenge for cross-turn coherence.
@@ 4.1.3 Ambiguity Recognition & Clarification
When inputs are ambiguous or underspecified, aligned LLMs tend to either over-hedge with vague answers or silently guess user intent rather than asking clarifying questions. Chen et al. (2025a) find that conversational agents frequently fail to disambiguate—responding generically rather than seeking clarification. An instruction such as "Tell me about that report" may elicit an arbitrary guess about which report rather than a targeted follow-up. Such unresolved ambiguities compound across turns, motivating research into LLMs that proactively recognize and resolve underspecification as humans naturally do.
@@ 4.2.1 Error Propagation & Compounding
Errors introduced in early turns—whether by the model or the user—tend to persist and amplify through subsequent dialogue. Once an incorrect answer or false premise enters the context, most LLMs continue building on it rather than self-correcting, leading to compounding reasoning failures (Kwan et al., 2024). This effect is especially pronounced in sequential decision tasks, where minor mistakes accumulate across interaction trajectories. The root difficulty is that models lack a robust mechanism to detect and retract mid-dialogue errors. Designing models that can identify mistakes and accept corrections—whether self-initiated or user-supplied—remains an open problem.